OpenStudy (anonymous):

What is the solution to 8w = 32? 0.25 4 24 40

OpenStudy (anonymous):

4

OpenStudy (anonymous):

divide both sides by 8, then you get the answer

OpenStudy (anonymous):

8w = 32 w = 32/8 w = 4

OpenStudy (anonymous):

use your eyeballs, get the answer. there is no difference between computing 32/8 and solving 8x = 32 both say "what times 8 is 32"?

OpenStudy (anonymous):

cool down satellite